# On the use of far-infrared and sub-millimeter line ratios to constrain the physical properties of gas in high-z galaxies

### v1

@@
On the use of far-infrared and sub-millimeter line ratios to constrain the physical properties of gas in high-z galaxies
Optical emission line diagnostics, which are a common tool to constrain the properties of the interstellar medium (ISM) of galaxies, become progressively inaccessible at higher redshifts for ground-based facilities. Far-infrared (FIR) emission lines, which are redshifted into atmospheric windows accessible by ground-based sub-millimeter facilities, could provide alternative ISM diagnostics to optical emission lines. We investigate FIR line ratios involving [CII] m, [OIII] m, [OIII] m, [NII] m and [NIII] m, using synthetic emission lines applied to a high-resolution (mgas= 883.4 M⊙) cosmological zoom-in simulation, including radiative-transfer post processing with Kramses-rt at z = 6.5. We find that the [CII]/[NII]122 ratio is sensitive to the temperature and density of photo-dissociation regions, and thus could be a useful tool to trace the properties of this gas phase in galaxies. We also find that [NII]/[NIII] is a good tracer of the temperature and [OIII]52/[OIII]88 a good tracer of the gas density of HII regions. Emission line ratios containing the [OIII] m line are sensitive to high velocity outflowing gas.
@@
Far-infrared (FIR) emission lines are useful tools for the investigation of the interstellar medium (ISM) of galaxies, as they are mostly unaffected by dust extinction (Brauher et al., 2008). In recent years, FIR emission lines have been used to explore the gas properties of galaxies such as ionisation parameter, gas density, gas metallicity, gas temperature and the radiation field (Nagao et al., 2011, 2013; Cormier et al., 2015; Pereira-Santaella et al., 2017; Rigopoulou et al., 2018; Killi et al., 2023; Ramos Padilla et al., 2023; Kumari et al., 2024). At several FIR lines are shifted into the sub-millimeter (sub-mm) atmospheric windows, where they can be observed by sensitive ground-based sub-mm telescopes, which makes them attractive tools to characterise the ISM of high-redshift galaxies in the early stages of their evolution.
In Schimek et al. (2024) we presented a line modelling effort including the [CII] m, [CI](1-0) m, [CI](2-1) m, CO(3-2) m, and [OIII] m lines. In this follow-up work, we adopt the same simulation and fiducial radiative transfer (RT) model as Schimek et al. (2024), and expand it to include the additional transitions of [OIII] m, [NII] m and [NIII] m, which trace the warm gas found in HII regions (Nagao et al., 2011). While in Schimek et al. (2024) the main subject was to study the extended sub-mm emission of the circum-galactic medium (CGM), we now focus on the inner part of the simulation. Inspired by analytical modelling of these line ratios as seen in Nagao et al. (2011) and Pereira-Santaella et al. (2017), we focus on studying how FIR emission line ratios change as functions of ISM properties when applied to a full galaxy simulation. Examples of previously explored FIR emission line ratios are the [OIII] m / [CII] m ratio, which has been investigated in both local and high-redshift galaxies (Vallini et al., 2015; Inoue et al., 2016; Carniani et al., 2018; Hashimoto et al., 2019; Pallottini et al., 2017, 2019; Katz et al., 2019; Harikane et al., 2020; Arata et al., 2020; Carniani et al., 2020; Lupi et al., 2020; Vallini et al., 2021; Fujimoto et al., 2024; Katz et al., 2022; Pallottini et al., 2022; Witstok et al., 2022; Kumari et al., 2024), and [OIII] m / [OIII] m, which has been shown to trace the gas density of ionised gas (Dinerstein et al., 1985; Rubin et al., 1994; Brauher et al., 2008; Nagao et al., 2011, 2013; Pereira-Santaella et al., 2017; Rigopoulou et al., 2018; Killi et al., 2023). Our results are aimed to inform observations performed with current and future sub-millimeter facilities such as the Atacama Large Millimeter/submillimeter Array (ALMA), the Atacama Pathfinder Experiment (APEX), the Northern Extended Millimeter Array (NOEMA) and, in the future, the Atacama Large Aperture Submillimetre Telescope (AtLAST)111https://www.atlast.uio.no (Mroczkowski et al., 2024).
@@ 2 Simulation and modelling
In this study we analyse the Ponos simulation (Fiacconi et al., 2017), a high resolution cosmological zoom-in simulation, run with the smoothed particle hydrodynamics (SPH) code Gasoline (Wadsley et al., 2004). The simulation has a gas mass resolution of m M⊙, a stellar particle mass of m m M⊙, and uses WMAP 7/9 cosmology (, , , , n, and H km s-1 Mpc-1) (Komatsu et al., 2011; Hinshaw et al., 2013). Ponos is the progenitor of a massive galaxy, undergoing a merger (stellar mass merger ratio 1 : 2.7) in the analysed snapshot at . The central galaxy has a stellar mass of M M⊙, a virial radius of R kpc, and a star formation rate SFR M∗ yr-1. These properties make Ponos a typical star forming galaxy at . Fig. 1 shows the gas density, temperature, metallicity and far-ultraviolet (FUV) radiation field of the main disc and the main merger of Ponos. Figures showing the entire halo can be seen in Schimek et al. (2024).
@@
We refer to Schimek et al. (2024) for the details of emission-line modelling, but briefly outline our fiducial model here. We first converted the Ponos simulation into an adaptive mesh refinement (AMR) grid, and then post-processed it with Kramses-rt (Pallottini et al., 2019; Decataldo et al., 2019, 2020), a customised version of Ramses-rt (Teyssier, 2002; Rosdahl et al., 2013) where a non-equilibrium chemical network generated via the package Krome (Grassi et al., 2014) has been implemented. We used the Kramses-rt accurate scheme for radiative transport, accounting for photo-reactions and gas self-shielding to compute the radiation in 10 bins for each grid cell. We computed stellar spectra for each star particle with Starburst 99 (SB99) (Leitherer et al., 1999), according to the age and the metallicity of the stellar particle, and then rescaled it to its mass. In addition to stellar radiation, we included a uniform ultraviolet background (UVB) (Haardt & Madau 2012 tables at ) and attenuated it in each cell according to the column density of each chemical species. The simulation was further post-processed with Cloudy (Ferland et al., 2017), creating a multi-dimensional grid for the line emission, based on the gas temperature, gas density, gas metallicity and the incident FUV radiation field. We include the radiation field of the gas with , which corresponds to the ionising FUV radiation field (6 - 13.6 eV) in units of Habing ( erg cm-2 s-1) (Habing, 1968).
@@
To compute line luminosity ratios and line ratio maps from the post-processed simulation data, we applied a luminosity thresholds of L L⊙ per fluid element, and L⊙ kpc-2 for the surface luminosity maps to both lines involved in the ratio. These thresholds are very low, and thus do not affect the integrated ratio values of the halo. The integrated line ratios for the entire halo and the different components can be seen in Table 1. We follow the same criteria to define the individual galaxy components as in Schimek et al. (2024), but in this study we focus on the disc and merger components within the central 25 kpc, since most lines are faint in the CGM. Due to the applied threshold, most of the far CGM is not included, as the CGM components that are bright enough to satisfy the criteria are located close to the main disc and merger components within the central 25 kpc. We classify the gas as ’inflows’ and ’outflows’, in the same way it was done in Schimek et al. (2024). We select the ‘high velocity’ gas with an absolute km s-1 (higher than the rotational velocity of the main disc), and define it as inflowing or outflowing by dividing it into gas that is flowing towards the main disc and away from the main disc.
@@ 3 Results
Our results are shown in Fig. 2, displaying emission line ratio maps, and Fig. 3, showing the ratios as a function of gas density, temperature, metallicity, incident FUV radiation field and absolute velocities of the gas cells. In the following, we comment on each line ratio separately.
@@ 3.1 O3C2
The [OIII] m / [CII] m (O3C2) ratio is of great interest, as the [CII] and the [OIII] emission lines are bright coolants of the ISM in the FIR, and can be observed with ground-based telescopes from a high-altitude site at and respectively. [CII] is one of the brightest FIR emission lines in star-forming (SF) galaxies, accounting for up to 0.1 - 1% of the total FIR emission (Stacey et al., 1991), and traces both ionised and neutral gas, due to the low excitation energy of carbon (11.26 eV, lower than hydrogen with 13.6 eV). [OIII] has a high ionisation energy of 35 eV, and traces hot and diffuse gas in HII regions and hot outflows. The O3C2 ratio has been investigated in both observational and theoretical studies of high- and low-redshift galaxies (Vallini et al., 2015; Inoue et al., 2016; Carniani et al., 2018; Hashimoto et al., 2019; Pallottini et al., 2017, 2019; Katz et al., 2019; Harikane et al., 2020; Arata et al., 2020; Carniani et al., 2020; Lupi et al., 2020; Vallini et al., 2021; Fujimoto et al., 2024; Katz et al., 2022; Pallottini et al., 2022; Witstok et al., 2022; Kumari et al., 2024; Ramos Padilla et al., 2023). As the [CII] and [OIII]88m lines trace different gas phases within the ISM, the analysis of the emission line ratio can give information about the physical conditions of the ISM, such as the PDR covering fraction, gas densities and temperatures.
@@
In Schimek et al. (2024) we showed that [CII] emission is more extended than [OIII], and while the respective line luminosities fall within the scatter of observational data, Ponos (total O3C2 ratio of 0.17) is more aligned with ratios in the local universe than with high-redshift galaxies which have observed ratios between 1 - 10 (Harikane et al., 2020). We note that local dwarf galaxies, often used as analogues of early galaxies, appear to share some of these extreme conditions, with [CII] being more extended than [OIII] (Cormier et al., 2015) and O3C2 ratios exceeding unity (Kumari et al., 2024). One of the reasons for the enhanced ratios measured at high-redshift could lie in abundance ratio differences due to an early stage of chemical evolution (see Nyhagen et al. in prep). The O3C2 map, reported in the upper left panel of Fig. 2, shows that the ratio is significantly enhanced in spatially compact areas in the disc, which correspond to hot, low density and high metallicity regions (see Fig.1).
@@
In Fig. 3, where the first row corresponds to the O3C2 ratio, we find that the O3C2 ratio is most sensitive to both gas density and temperature, as it varies by four orders of magnitude, being low in high density (n cm-3) and low temperature gas (T K), and high in low density (n cm-3) and high temperature gas (T K). We find O3C2 ratio values of 1 for typical values of HII regions (T 7103 - 1.5104 K, n 10 - 102 cm-3, Osterbrock & Ferland (2006)). The highest ratios (10) are found in gas with typical densities and temperatures of SNe remnants (T 105.7 K, n 10-2.5 cm-3, McKee & Ostriker (1977)). We see a small scatter for the ratio as a function of the density and temperature, and also do not see significant differences between the ISM components of the disc and the merger and the CGM, probably because this analysis, as explained in Section2, is not sensitive to the most extended CGM gas. The dependence of O3C2 on the radiation field seems not very significant up to log10(G0) 2, above which the ratio increases as the strong radiation can ionise [OIII], while carbon is doubly ionised, decreasing the [CII] emission. As we assume uniform solar abundance ratios for the gas in our Cloudy modelling, which are scaled with the metallicity, the lack of dependence of the O3C2 ratio on the gas phase metallicity for sub-solar metallicities is not surprising. For all metallicities, the C/O ratio is constant in our models, while in galaxies the C/O ratio would be metallicity dependent due to the different timescales of enrichment processes (Maiolino & Mannucci, 2019; Arata et al., 2020). At super-solar metallicities, we see a sharp increase of the O3C2 ratio by two orders of magnitude. We note that only a small fraction of the gas mass in Ponos (5.2%) reaches such high metallicities. The increase of the ratio as a function of metallicity is caused by the properties of the gas, as it seems to have been recently enriched by core-collapse supernova (cc-SNe), and is generally heated, has low densities, and experiences a strong radiation field. At log10(Z/Z⊙) 0.2 we find gas with O3C2 10, with temperatures of 104 - 2104 K, densities of n 0.1 - 10 cm-3. It also experiences strong radiation fields of log(G0) 2.0 - 2.5. We attribute the sharp increase of the O3C2 emission line ratio for super-solar metallicities to the effect of the radiation field, as well as the low gas densities and high temperatures. From observational results (Cormier et al., 2015) and theoretical modelling where individual elements are traces (Arata et al., 2020) we would expect lower O3C2 ratios at higher metallicities. Directly tracing carbon and oxygen with their respective yields caused by different enrichment events (e.g. cc-SNe, SN Type Ia, asymptotic giant branch (AGB) stars, etc.) could reveal a significant metallicity dependence. Arata et al. (2020) found in their semi-analytical modelling that the O3C2 ratio decreases by two orders of magnitude with increasing metallicity, which could counteract the trend we find when assuming a fixed, metallicity independent C/O ratio. The ratio has a larger scatter as a function of both gas phase metallicity and radiation field, compared to density and temperature ( 0.5 dex). The last column of Fig. 3 shows the ratio as a function of the absolute velocity of the gas, divided into inflowing and outflowing components. It can be seen that outflowing gas shows higher O3C2 ratios compared to inflowing gas, due to gas heated by SNe feedback, where [OIII] radiation dominates.

### v2

# Constraining the physical properties of gas in high-z galaxies with far-infrared and submillimetre line ratios

@@
Constraining the physical properties of gas in high-z galaxies with far-infrared and submillimetre line ratios
Optical emission line diagnostics, which are a common tool for constraining the properties of the interstellar medium (ISM) of galaxies, become progressively inaccessible at higher redshifts for ground-based facilities. Far-infrared (FIR) emission lines, which are redshifted into atmospheric windows that are accessible for ground-based submillimetre facilities, could provide ISM diagnostics alternative to optical emission lines. We investigated FIR line ratios involving [CII] m, [OIII] m, [OIII] m, [NII] m, and [NIII] m using synthetic emission lines applied to a high-resolution (mgas= 883.4 M⊙) cosmological zoom-in simulation, including radiative transfer post-processing with the code Kramses-rt at z = 6.5. We find that the [CII]/[NII]122 ratio is sensitive to the temperature and density of photodissociation regions. It might therefore be a useful tool for tracing the properties of this gas phase in galaxies. We also find that [NII]/[NIII] is a good tracer of the temperature and that [OIII]52/[OIII]88 is a good tracer of the gas density of HII regions. Emission line ratios containing the [OIII] m line are sensitive to high-velocity outflowing gas.
@@
Far-infrared (FIR) emission lines are useful tools for the investigation of the interstellar medium (ISM) of galaxies because they are mostly unaffected by dust extinction (Brauher et al., 2008). In recent years, FIR emission lines have been used to explore the gas properties of galaxies, such as the ionisation parameter, gas density, gas metallicity, gas temperature, and the radiation field (Nagao et al., 2011, 2013; Cormier et al., 2015; Pereira-Santaella et al., 2017; Rigopoulou et al., 2018; Killi et al., 2023; Ramos Padilla et al., 2023; Kumari et al., 2024). At , several FIR lines are shifted into submillimetre (submm) atmospheric windows, where they can be observed by sensitive ground-based submm telescopes. This makes them attractive tools for characterising the ISM of high-redshift galaxies in the early stages of their evolution.
In Schimek et al. (2024) we presented a line-modelling effort that included the [CII] m, [CI](1-0) m, [CI](2-1) m, CO(3-2) m, and [OIII] m lines. In this follow-up work, we adopt the same simulation and fiducial radiative transfer model as Schimek et al. (2024) and expand it to include the additional transitions of [OIII] m, [NII] m, and [NIII] m, which trace the warm gas found in HII regions (Nagao et al., 2011). While in Schimek et al. (2024) the main subject was to study the extended submm emission of the circumgalactic medium (CGM), we now focus on the inner part of the simulation. Inspired by analytical modelling of these line ratios as seen in Nagao et al. (2011) and Pereira-Santaella et al. (2017), we focus on the change in the FIR emission line ratios as functions of ISM properties when applied to a full galaxy simulation. Examples of previously explored FIR emission line ratios are the [OIII] m / [CII] m ratio, which has been investigated in both local and high-redshift galaxies (Vallini et al., 2015; Inoue et al., 2016; Carniani et al., 2018; Hashimoto et al., 2019; Pallottini et al., 2017, 2019; Katz et al., 2019; Harikane et al., 2020; Arata et al., 2020; Carniani et al., 2020; Lupi et al., 2020; Vallini et al., 2021; Fujimoto et al., 2022; Katz et al., 2022; Pallottini et al., 2022; Witstok et al., 2022; Kumari et al., 2023), and [OIII] m / [OIII] m, which has been shown to trace the gas density of ionised gas (Dinerstein et al., 1985; Rubin et al., 1994; Brauher et al., 2008; Nagao et al., 2011, 2013; Pereira-Santaella et al., 2017; Rigopoulou et al., 2018; Killi et al., 2023). Our results are aimed to inform observations performed with current and future submm facilities, such as the Atacama Large Millimeter/submillimeter Array (ALMA), the Atacama Pathfinder Experiment (APEX), the Northern Extended Millimeter Array (NOEMA), and, in the future, the Atacama Large Aperture Submillimeter Telescope (AtLAST)111https://www.atlast.uio.no (Mroczkowski et al., 2024).
@@ 2 Simulation and modelling
In this study, we analyse the Ponos simulation (Fiacconi et al., 2017), a high-resolution cosmological zoom-in simulation that was run with the smoothed particle hydrodynamics (SPH) code Gasoline (Wadsley et al., 2004). The simulation has a gas mass resolution of m M⊙ and a stellar particle mass of m m M⊙, and it uses the WMAP 7/9 cosmology (, , , , n, and H km s-1 Mpc-1) (Komatsu et al., 2011; Hinshaw et al., 2013). Ponos is the progenitor of a massive galaxy, undergoing a merger (stellar mass merger ratio 1 : 2.7) in the analysed snapshot at . The central galaxy has a stellar mass of M M⊙, a virial radius of R kpc, and a star formation rate, SFR, of M∗ yr-1. These properties make Ponos a typical star-forming galaxy at . Fig. 1 shows the gas density, temperature, metallicity, and far-ultraviolet (FUV) radiation field of the main disc and the main merger of Ponos. Figures showing the entire halo were presented in Schimek et al. (2024).
@@
We refer to Schimek et al. (2024) for the details of emission-line modelling, but briefly outline our fiducial model here. We first converted the Ponos simulation into an adaptive mesh refinement (AMR) grid and then post-processed it with Kramses-rt (Pallottini et al., 2019; Decataldo et al., 2019, 2020), a customised version of Ramses-rt (Teyssier, 2002; Rosdahl et al., 2013), where a non-equilibrium chemical network generated via the package Krome (Grassi et al., 2014) has been implemented. We used the Kramses-rt accurate scheme for radiative transport, accounting for photoreactions and gas self-shielding to compute the radiation in ten bins for each grid cell. We computed stellar spectra for each stellar particle with Starburst 99 (SB99) (Leitherer et al., 1999), according to the age and the metallicity of the stellar particle, and we then rescaled it to its mass. In addition to stellar radiation, we included a uniform ultraviolet background (UVB) (Haardt & Madau 2012 tables at ) and attenuated it in each cell according to the column density of each chemical species. The simulation was further post-processed with Cloudy (Ferland et al., 2017), creating a multi-dimensional grid for the line emission, based on the gas temperature, gas density, gas metallicity, and the incident FUV radiation field. We included the radiation field of the gas with , which corresponds to the ionising FUV radiation field (6 - 13.6 eV) in units of Habing ( erg cm-2 s-1) (Habing, 1968).
@@
To compute line luminosity ratios and line ratio maps from the post-processed simulation data, we applied a luminosity thresholds of L L⊙ per fluid element and L⊙ kpc-2 for the surface luminosity maps to the two lines involved in the ratio. These thresholds are very low and therefore do not affect the integrated ratio values of the halo. The integrated line ratios for the entire halo and the different components can be seen in Table 1. We followed the same criteria to define the individual galaxy components as in Schimek et al. (2024), but in this study, we focus on the disc and merger components within the central 25 kpc because most lines in the CGM are faint. The applied threshold means that most of the far CGM is not included because the CGM components that are bright enough to satisfy the criteria are located close to the main disc and merger components within the central 25 kpc. We classified the gas into ’inflows’ and ’outflows’, in the same way as in Schimek et al. (2024). We selected ‘high-velocity’ gas with an absolute km s-1 (higher than the rotational velocity of the main disc), and defined it as inflowing or outflowing by dividing it into gas that flowed towards and away from the main disc, respectively.
@@ 3 Results
Our results are shown in Fig. 2, displaying emission line ratio maps, and Fig. 3, showing the ratios as a function of gas density, temperature, metallicity, incident FUV radiation field, and absolute velocities of the gas cells. In the following, we comment on each line ratio separately.
@@ 3.1 O3C2
The [OIII] m / [CII] m (O3C2) ratio is of great interest because the [CII] and the [OIII] emission lines are bright coolants of the ISM in the FIR and can be observed with ground-based telescopes from a high-altitude site at and , respectively. [CII] is one of the brightest FIR emission lines in star-forming (SF) galaxies. It accounts for up to 0.1 - 1% of the total FIR emission (Stacey et al., 1991) and traces both ionised and neutral gas because of the low excitation energy of carbon (11.26 eV, which is lower than hydrogen with 13.6 eV). [OIII] has a high ionisation energy of 35 eV and traces hot and diffuse gas in HII regions and hot outflows. The O3C2 ratio has been investigated in both observational and theoretical studies of high- and low-redshift galaxies (Vallini et al., 2015; Inoue et al., 2016; Carniani et al., 2018; Hashimoto et al., 2019; Pallottini et al., 2017, 2019; Katz et al., 2019; Harikane et al., 2020; Arata et al., 2020; Carniani et al., 2020; Lupi et al., 2020; Vallini et al., 2021; Fujimoto et al., 2022; Katz et al., 2022; Pallottini et al., 2022; Witstok et al., 2022; Kumari et al., 2023; Ramos Padilla et al., 2023). As the [CII] and [OIII]88m lines trace different gas phases within the ISM, the analysis of the emission line ratio can give information about the physical conditions of the ISM, such as the PDR covering fraction, gas densities, and temperatures.
@@
In Schimek et al. (2024) we showed that [CII] emission is more extended than [OIII], and while the respective line luminosities fall within the scatter of observational data, Ponos (total O3C2 ratio of 0.17) is more aligned with ratios in the local Universe than with high-redshift galaxies, which have observed ratios between 1 - 10 (Harikane et al., 2020). We note that local dwarf galaxies, which are often used as analogues of early galaxies, appear to share some of these extreme conditions, with [CII] being more extended than [OIII] (Cormier et al., 2015) and O3C2 ratios exceeding unity (Kumari et al., 2024). One of the reasons for the enhanced ratios measured at high-redshift could lie in abundance ratio differences that are due to an early stage of chemical evolution (see Nyhagen et al. in prep). The O3C2 map reported in the upper left panel of Fig. 2 shows that the ratio is significantly enhanced in spatially compact areas in the disc, which correspond to hot low-density and high-metallicity regions (see Fig.1).
@@
In Fig. 3, where the first row corresponds to the O3C2 ratio, we find that the O3C2 ratio is most sensitive to the gas density and temperature because it varies by four orders of magnitude. It is low in high-density (n cm-3) and low-temperature gas (T K) and high in low-density (n cm-3) and high-temperature gas (T K). We find O3C2 ratio values of 1 for typical values of HII regions (T 7103 - 1.5104 K, n 10 - 102 cm-3, Osterbrock & Ferland (2006)). The highest ratios (10) are found in gas with typical densities and temperatures of SNe remnants (T 105.7 K, n 10-2.5 cm-3; McKee & Ostriker (1977)). The ratio shows little scatter as a function of the density and temperature, and the ISM components of the disc and the merger and the CGM are similar, probably because this analysis, as explained in Section2, is not sensitive to the most extended CGM gas. The dependence of O3C2 on the radiation field seems not very significant up to log10(G0) 2, above which the ratio increases as the strong radiation can ionise [OIII], while carbon is doubly ionised, which decreases the [CII] emission. As we assumed uniform solar abundance ratios for the gas in our Cloudy modelling, which are scaled with the metallicity, the lack of a dependence of the O3C2 ratio on the gas-phase metallicity for subsolar metallicities is expected. For all metallicities, the C/O ratio is constant in our models, while in galaxies, the C/O ratio would depend on metallicity because the timescales of the enrichment processes are different (Maiolino & Mannucci, 2019; Arata et al., 2020). At supersolar metallicities, the O3C2 ratio increases sharply by two orders of magnitude. We note that only a small fraction of the gas mass in Ponos (5.2%) reaches high metallicities like this. The increase in the ratio as a function of metallicity is caused by the properties of the gas, as it seems to have recently been enriched by core-collapse supernova (cc-SNe) and is generally heated, has low densities, and experiences a strong radiation field. At log10(Z/Z⊙) 0.2, we find gas with O3C2 10, with temperatures of 104 - 2104 K and densities of n 0.1 - 10 cm-3. It also experiences strong radiation fields of log(G0) 2.0 - 2.5. We attribute the sharp increase in the O3C2 emission line ratio for supersolar metallicities to the effect of the radiation field and to the low gas densities and high temperatures. From observational results (Cormier et al., 2015) and theoretical modelling, where individual elements are traced (Arata et al., 2020), we would expect lower O3C2 ratios at higher metallicities. A direct tracing of carbon and oxygen with their respective yields caused by different enrichment events (e.g. cc-SNe, SN Type Ia, and asymptotic giant branch (AGB) stars) could reveal a significant metallicity dependence. Arata et al. (2020) found in their semi-analytical modelling that the O3C2 ratio decreases by two orders of magnitude with increasing metallicity, which could counteract the trend we find when we assumed a fixed metallicity-independent C/O ratio. The ratio has a larger scatter as a function of both gas-phase metallicity and radiation field compared to density and temperature ( 0.5 dex). The last column of Fig. 3 shows the ratio as a function of the absolute velocity of the gas, divided into inflowing and outflowing components. The outflowing gas shows higher O3C2 ratios than the inflowing gas, which is due to gas heated by SNe feedback, where [OIII] radiation dominates.